# Empirically Exploring How Novices Write Software Models in Alloy

### v2

@@ 4.4. RQ4: How do novice users’ correct answers differ from the oracle formula?
Often times, the oracle solution represents one of several ways in which the property could be expressed, as seen in Tables 3 and 4 in which the number of syntactically unique correct solutions is larger than the number of semantically unique correct solutions. Therefore, we wanted to explore what ways novice users’ thinking may differ from the expert written oracles.
First, we explore how frequently the oracle solution appears as a correct submission. Table 6 displays the location of the oracle submission derived by ranking all syntactically unique correct submissions in decreasing order from most to least duplicate submissions. Then, we looked for the location of the oracle submission within that list. Column Type displays the type of logical exercises under consideration based on the oracle formula. Column Top 1 shows the number of exercises for that type in which the oracle is first in the ranked list and associated Column Rate is what percentage this number is out of the total number of exercises of that type. Table 6 also displays information for whether the oracle is in the Top 5, Top 10 or never submitted (Not Sub).
@@
For RL exercises, the oracle is the top submission 43.14% of the times and in the top 10 96.08% of the time. Rarely (5.33%) the oracle is never submitted. The oracle is only the top submission for 26.32% of FOL exercises and 33.33% of LTL exercises. By location 10, the oracle is only present in 68.42% and 61.11% of the exercises respectively. Moreover, it is common for the oracle to never be submitted, which happens about 25% of the time for both FOL and LTL exercises. RL exercise use the smallest subset of operators, and this seems to translate to less diversity in how a property can be correctly expressed. This is further supported by Table 3, in which 15.7% of correct RL submissions are syntactically unique but 21.85% of FOL submissions are syntactically unique. Interestingly, only 12.58% of LTL exercises are syntactically unique, which indicates that users frequently write the same correct submissions, but fail to utilize the same structure as the oracle.
@@
Nearly one third of the top 1 correct submissions are syntactically equivalent to the oracle. For the remaining two-thirds, 35 submissions are similar, but involve minor differences, such swapping the order of operands for commutative operators. Many of these submissions are LTL exercises in which the location of the temporal operator always varies. In addition, the three of the four “similar but additional trivial operator” are LTL exercises in which the user appends an unnecessary leading always operator. Together, these similar submissions highlight common modeling preferences that result in the oracle being less present for LTL exercises.
@@
Overall, Table 8 reinforces the implication from Table 7 that novice users often write larger formulas. RL and PL exercises see the largest increase in size, which is expected, as one of the common reasons the formula is larger based on Table 7 is that the user is adding at least one new quantified formula to their solution compared to the oracle. Not surprisingly, this difference in the number of AST nodes translates to the oracle being on average 0.38$\times$ smaller for both RL and PL exercises. However, even FOL exercises, which already have a quantified formula, also see a large increase in the number of AST nodes, with the oracle on average being 8.17 nodes smaller. Table 7 highlights that users even up-scaled existing quantified formulas to have nested or even further nested quantification. For LTL exercises, while the novice correct submissions are still on average larger, the difference is only 3.7 nodes. However, these 3.7 nodes results in an oracle that is on average 0.77$\times$ the size of the correct submission, which is similar to FOL exercises in which the 8.17 nodes means the oracle is on average 0.81$\times$ the size of the correct submissions.
@@
[…]
@@ 4.5. RQ5: How often do novice users reach a correct answer?
Considering that 25.4% of the submissions are correct, we were curious to explore how novice users arrive at the correct answer. For each model, Table 9 displays what percentage of attempts have a correct submission, what percentage of attempts have a correct answer as the first submission, what percentage of attempts never produce a correct answer, and for those without a correct answer, the average number of submissions made until the user gave up. Since the user may switch between different types of logic within each attempt, we do not report any type-based information.
Across all attempts, 67.38% of the times the user gets to a correct answer and notably users write a correct answer first in 31.23% of all attempts. The complexity of the underlying base model, as captured in Table 1, does have some roll in users reaching a correct submission, as the trash models and graph models have some of the highest rate of attempts with a correct submission. However, the complexity of the underlying base model does not necessarily translate to how often novices get solutions correct on the first attempt. For instance, while trash_fol has the highest percentage and is a simple model, classroom_fol has the second highest percentage and is notable more complex. Instead, correct on first attempts were more closely connected to exercises that occurred earlier in a model. Most models are set up such that the difficulty of the exercise increases as you progress through the exercises.
@@
Novice users struggle with courses_v1, courses_v2 and cv_v1, in which less than 50% of all attempts made for these models have a correct submission. Both courses_v1 and courses_v2 have two signatures that contain a relation with the same name, which adds additional complexity to the exercises as users often have to properly scope formulas to remove ambiguity in which relation is getting accessed. In addition, four models with low correctness rates, cv_v1, cv_v2, socialMedia and trainStation_ltl, are all models that require first order logic to solve most exercises. Furthermore, socialMedia and trainStation_ltl have the largest oracle submissions in terms of number of AST nodes, indicating that the properties-to-be-expressed for these models are more complex compared to the other models. Complexity of the exercise likely influences more than just how often users reach a correct solution. Across all models, there is a dropoff in the number of overall attempts for the later exercises for a model. This indicates that users tend to stop working on a model as the exercises get harder.
@@ 4.6. RQ6: How often do novice users make the same logical mistake?
As indicated by Table 3, novice users often repeat the same exact mistake. Understanding these common mistakes can help researchers focusing on fault localization and repair, as well as educators teaching formal methods. To give further guidance on this, we explore a subset of “hot hit” submissions in which the same syntactic submission is made at least 10 times. We manually investigated each “hot hit” and label it with core tenants of the mistake. Table 10 displays the common mistakes made broken down by classification of submission and type of logic. Column #Occ represent the number of unique “hot hit” submissions of that mistake and column #Sub represents how many total number of all “hot hit” submissions of that mistake. In green, we highlight the top 4 mistakes.
Across all logic types, the single most common mistake is the incorrect use of an operator. Our dataset on Github includes a detailed breakdown of which specific operators are incorrectly applied and at what rates. The most common misapplication is related to the use of the subset, relational join and implication operators. For subset mistakes, a common but very subtle mistake is for novice users to use subset in place of set intersection. For their formulas, these two operators often behave the same except for when the left-hand side of the operator is empty. For example, consider the following submission (s) to inv6 from the productionLine_v2 model that 20 novice users submitted (19.61% of inv7’s both over- and underconstrained submissions) and the closest correct solution (c):
@@
With FOL submissions prevalent in Table 10, there are fundamental misunderstandings novice users have about the appropriate application of FOL. Half of the frequently repeated classification of mistakes (column Mistakes) involve novices introducing mistakes related to quantification, which reiterates our observations in RQ1 that FOL formulas gave novice users problems. The most common quantifier mistake, and second most overall mistake, is that novices try to upscale the level of quantification in an attempt to explicitly outline relationships, where upscale means the submission inserted more quantifiers than the oracle. As an example, consider the following submission for classroom_rl inv5:
@@
that accounts for 30 (35.29%) of the both over- and underconstrained submissions for inv5. To correct this formula, the user should change the quantification universal quantification (‘all’) to existential quantification (‘some’). While there are often ways for to express a relational formula with an equivalent first order logic formula, the increase in complexity of the formula structure introduces more opportunities for users to introduce a mistake.
Hand in hand with upscaling, downscaling the level of quantification is a common mistake that often results in an expressions being applied to the wrong set of elements. As an example, consider the submission for inv1 for model lts:
@@
The forth most common overall mistake is for users to make the domain of a quantified formula too narrow, which frequently produces an underconstrained submission. A narrower quantification domain almost always occurred on models that had extension signatures and the novice user used an extension signature for the domain in place of the base signature. As an example, consider the following submission for inv3 for model courses_v2 which accounts for 65 (40.37%) of the underconstrained submissions:
@@
These narrowed quantification domains additional account for the third most common mistake, the incorrect use of an extension signature. In addition to incorrectly using extension signatures in quantified domains, novice users also incorrectly used extension signatures within individual relational expressions.
@@
In addition, these common mistakes provide clear blueprints for researchers to aid in improving debugging-oriented techniques.
@@
[…]
@@ 4.7. RQ7: How often are users “one” mistake away?
For incorrect submissions, we were curious to explore how often users were effectively “one mistake away” from a correct solution. To investigate this, we applied MuAlloy, a mutation testing tool (MuAlloyTemporal, ), to all faulty submissions in the both, over and under categories. For each submission, we generated all first order mutants and then checked if any of the mutated models were equivalent to the oracle submission. If no counterexample is found, we classified the submission as fixable and “one mistake away.”
[…]
Overall, only 10.86% of submissions can be correct by the current first order mutants operators. The rate is notable lowest for RL submissions, in which mutants can only correct 5.61%. In contrast, LTL submission have nearly quadruple the fix rate at 19.74%. This trend holds for submissions with multiple fixes as well. Based on classification, overconstrained faults have the most fixes possible at 18.34% followed by underconstrained faults at 15.48%, while both over- and underconstrained faults can only be fixed 6.24% of the time. This is expected, as both over- and underconstrained faults mean the formula is sometimes too permissive and sometimes too restrictive, which is likely to lead to multi-step edits not captured by first order mutants.
The overall low fix rate is in line with observations from RQ6 in which users over make very similar logical mistakes but the fix is actually two edit steps away, not one. For instance, there is no first order mutation operator that transforms “a in b,” into “some a & b.” In addition, there is no first order mutant operator to transform “all a : A | some a.b” into “some A.b.” while second order mutants would have scalability issues, the common mistakes from RQ6 could be utilized to produce a select subset that reflect common mistakes. Interestingly, mutation operators can correct some type errors, such as transforming “s.enrolled.grades” from RQ4 into “some s.enrolled.grades,” which could be utilized to improve the Analyzer’s current type-based error reports.
@@
The high rate of replacement fixes is supported by observations from our common mistakes (RQ6), where we found novice users incorrectly selecting between similar but different operators. In particular, BOR has the second highest fix rate (19.89%), which is reflective of the fact that a vast majority of the operators that were mixed up are binary operators. QOR fixes having the third highest fix rate (17.78%) is also reflective of both choosing the wrong operator, as well as the high prevalence of FOL mistakes. Interestingly, we found that several of the QOR fixes are related to replacing one of the nested quantifiers, such as replacing “all a, some b” with “all a, one b”. For insertions, UOI fixes were dominate at 31.16% of all fixes. UOI fixes for LTL operators almost always involve inserting always, which is not unexpected as RQ4 highlights that novice users tend to push temporal operators to the front of formulas. For RL, PL and FOL submissions, the UOI fixes were often the result of inserting transitive closure or transpose, the later of which is interesting as it means the user was close to the write express but had inverted part of the logic.
@@ 4.8. RQ8: How do novice users respond to faulty or invalid submissions?
As users incrementally work towards a correct solution, users are likely to make changes that may move their submission from one classification to the next. These trends can provide insight into how users respond to discovering their submission is faulty or invalid. To explore this, we collected all back to back single edit steps for an attempt. Table 13 displays the classification change, or lack there of, produced by all of these edits. The letter E represents empty submissions. Row # Edit ‘?’ represents the number of edits that started as one classification and moved to a new one based on the corresponding column, while Row % Edit ‘?’ represents what percentage this number is out of the total edits that start as classification ‘?’. For instance there are 4,133 edits that in which the user starts with an overconstrained formula and produces an overconstrained formula (O-O), which accounts for 40.5% of the edits that start with an overconstrained formula.
Based on Table 13, a user is significantly more likely to make an edit that does not change the classification, which implies that users often make small edits. In addition, users do frequently go from an overconstrained to a correct formula (22.92%), which is in line with our “one step away” results in which overconstrained formulas were often close to a correct submission. Interestingly, users do go from a correct submission to all the other classifications. There are two main reasons for this. First, some of the exercise ask for an answer using a specific type of logic, but Alloy4Fun does not restrict the available operators. Therefore, users will create a correct RL submission and attempt to re-write it into a FOL submission. Second, users would refine their submission, trying to condense the formula or try to produce more readable formulas.

### v3

@@ 4.4. RQ4: How do novice users’ correct answers differ from the oracle formula?
Oftentimes, the oracle solution represents one of several ways in which the property could be expressed, as seen in Tables 3 and 4 in which the number of syntactically unique correct solutions is larger than the number of semantically unique correct solutions. Therefore, we wanted to explore the ways novice users’ thinking may differ from the expert-written oracles.
First, we explore how frequently the oracle solution appears as a correct submission. Table 6 displays the location of the oracle submission derived by ranking all syntactically unique correct submissions in decreasing order from most to least duplicate submissions. Then, we looked for the location of the oracle submission within that list. Column Type displays the type of logical exercises under consideration based on the oracle formula. Column Top 1 shows the number of exercises for that type in which the oracle is first in the ranked list and associated Column Rate is what percentage this number is out of the total number of exercises of that type. Table 6 also displays information for whether the oracle is in the Top 5, Top 10, or never submitted (Not Sub).
@@
For RL exercises, the oracle is the top submission 43.14% of the time and in the top 10 96.08% of the time. Rarely (5.88%) the oracle is never submitted. The oracle is only the top submission for 26.32% of FOL exercises and 33.33% of LTL exercises. By location 10, the oracle is only present in 68.42% and 61.11% of the exercises respectively. Moreover, it is common for the oracle to never be submitted, which happens about 25% of the time for both FOL and LTL exercises. RL exercise use the smallest subset of operators, and this seems to translate to less diversity in how a property can be correctly expressed. This is further supported by Table 3, in which 15.7% of correct RL submissions are syntactically unique but 21.85% of FOL submissions are syntactically unique. Interestingly, only 12.58% of LTL exercises are syntactically unique, which indicates that users frequently write the same correct submissions, but fail to utilize the same structure as the oracle.
@@
Nearly one-third of the top 1 correct submissions are syntactically equivalent to the oracle. For the remaining two-thirds, 35 submissions are similar but involve minor differences, such swapping the order of operands for commutative operators. Many of these submissions are LTL exercises in which the location of the temporal operator always varies. In addition, the three of the four “similar but additional trivial operator” are LTL exercises in which the user appends an unnecessary leading always operator. Together, these similar submissions highlight common modeling preferences that result in the oracle being less present for LTL exercises.
@@
Table 8 displays the Abstract Syntax Tree (AST) difference between the oracle solution and all of the correct submissions. Column Type conveys the type of logic, with a final row that reflects all correct submission. Column DifTot displays the average difference in total number of AST nodes between the oracle and all correct submissions. The next four columns (DifRL, DifPL, DifFOL, DifLTL) displays the average difference in total number of AST nodes of each type of logic (RL, PL, FOL, and LTL) between the oracle and all correct submissions. For these columns, a negative number means the oracle had less AST nodes. To give context to how significant the difference in nodes is, column DifMag displays average magnitude in difference between the oracle and correct submissions.
@@
DifTot DifRL DifPL DifFOL DifLTL DifMag
@@ 4.5. RQ5: How often do novice users reach a correct answer?
Since only a forth of all submissions are correct, there is a high likelihood that some attempts end without reaching a correct answer. To explore this, for each model, Table 9 displays what percentage of attempts have a correct submission, what percentage of attempts have a correct answer as the first submission, what percentage of attempts never produce a correct answer, and for those without a correct answer, the average number of submissions made until the user gave up. Since the user may switch between different types of logic within each attempt, we do not report type-based information.
Across all attempts, 67.38% of the time the user gets to a correct answer and notably users write a correct answer first in 31.23% of all attempts. The complexity of the underlying base model, as captured in Table 1, does have some roll in users reaching a correct submission, as the trash models and graph models have some of the highest rate of attempts with a correct submission. However, the complexity of the underlying base model does not necessarily translate to how often novices get solutions correct on the first attempt. For instance, while trash_fol has the highest percentage and is a simple model, classroom_fol has the second highest percentage and is notable more complex. Instead, correct on first attempts were more closely connected to exercises that occurred earlier in a model. Most models are set up such that the difficulty of the exercise increases as you progress through the exercises.
@@
Novice users struggle with courses_v1, courses_v2 and cv_v1, in which less than 50% of all attempts made for these models have a correct submission. Both courses_v1 and courses_v2 have two signatures that contain a relation with the same name, which adds additional complexity to the exercises as users often have to properly scope formulas to remove ambiguity in which relation is getting accessed. For cv_v1 and additionally three other models low correctness rates – cv_v2, socialMedia and trainStation_ltl – all of these models require first order logic to solve most of their exercises. Furthermore, socialMedia and trainStation_ltl have the largest oracle submissions in terms of number of AST nodes, indicating that the properties-to-be-expressed for these models are more complex compared to the other models. Complexity of the exercise likely influences more than just how often users reach a correct solution. Across all models, there is a dropoff in the number of attempts for the later exercises for a model. This indicates that users tend to stop working on a model as the exercises get harder.
@@ 4.6. RQ6: How often do novice users make the same logical mistake?
As indicated by Table 3, novice users often repeat the same mistake. Understanding these common mistakes can help researchers focusing on fault localization and repair, as well as educators teaching formal methods. To give further guidance on this, we explore a subset of “hot hit” submissions in which the same syntactic submission is made at least 10 times. We manually investigated each “hot hit” and label it with core tenants of the mistake. Table 10 displays the common mistakes made broken down by classification of submission and type of logic. Column #Occ represent the number of unique “hot hit” submissions of that mistake and column #Sub represents how many total number of all “hot hit” submissions of that mistake. In green, we highlight the top 4 mistakes.
Across all logic types, the single most common mistake is the incorrect use of an operator. Our dataset includes a detailed breakdown of which specific operators are incorrectly applied and at what rates. The most common misapplication is related to the use of the subset, relational join, and implication operators. For subset mistakes, a common but very subtle mistake is for novice users to use subset in place of set intersection. For their formulas, these two operators often behave the same except for when the left-hand side of the operator is empty. For example, consider the following submission (s) to inv6 from the productionLine_v2 model that 20 novice users submitted (19.61% of inv7’s both over- and underconstrained submissions) and the closest correct solution (c):
@@
FOL submissions are prevalent in Table 10, accounting for 468 of the 565 hot hit submissions. In addition, half of the frequently repeated classification of mistakes (column Mistakes) involve novices introducing errors related to quantification. Therefore, there are clearly fundamental misunderstandings novice users have about how and when to use quantification. The most common quantifier mistake, and the second most common overall mistake, is that novices try to upscale the level of quantification in an attempt to explicitly outline relationships, where upscale means the submission inserted more quantifiers than the oracle. As an example, consider the following submission for classroom_rl inv5:
@@
that accounts for 30 (35.29%) of the both over- and underconstrained submissions for inv5. To correct this formula, the user should exchange universal quantification (‘all’) to existential quantification (‘some’). While there are often ways to express a relational formula with an equivalent first-order logic formula, the increase in complexity of the formula structure introduces more opportunities for users to make a mistake.
Hand in hand with upscaling, downscaling the level of quantification is a common mistake that often results in an expression being applied to the wrong set of elements. As an example, consider the submission for inv1 for model lts:
@@
The fourth most common overall mistake is for users to make the domain of a quantified formula too narrow, which frequently produces an underconstrained submission. A narrower quantification domain almost always occurred on models that had extension signatures and the novice user used an extension signature for the domain in place of the base signature. As an example, consider the following submission for inv3 for model courses_v2 which accounts for 65 (40.37%) of the underconstrained submissions:
@@
These narrowed quantification domains additionally account for the third most common mistake, the incorrect use of an extension signature. In addition to incorrectly using extension signatures in quantified domains, novice users also incorrectly used extension signatures within individual relational expressions.
@@
Hand in hand with education, these guidelines are also blueprints for researchers to aid in improving debugging-oriented techniques, such as creating new repair templates or refining suspiciousness formulas for fault localization.
@@
Clx/Type # Fixes # Subf % Subf # Subm % Subm
@@ 4.7. RQ7: How often are users “one” mistake away?
For incorrect submissions, we were curious to explore how often users were effectively “one mistake away” from a correct solution. To investigate this, we applied MuAlloy, a mutation testing tool (MuAlloyTemporal, ), to all faulty submissions in the both, over and under categories. For each submission, we generated all first-order mutants. Then, if any of the mutated models are equivalent to the oracle submission we classify the submission as fixable and “one mistake away.”
We were interested in two viewpoints: (1) how often are submissions one mistake away and (2) what mutants correct faulty submissions. For the first viewpoint, Table 11 depicts the frequency at which mutants fix submissions broken down by the type of logic present in the submission and the classification of the submission. Column #Fixes displays the total number of mutants that fixed a submission, column # Subf shows the number of submissions that were fixed and column % Subf shows what percentage of submissions # Subf is. The number of fixes is more than the number of submissions fixed because some submissions could be fixed by multiple different mutations. To help explore this, column # Subm shows the number of submissions that have multiple fixes, and column % Subm shows what percentage of submissions # Subm is.
Overall, only 10.86% of submissions can be corrected by the current mutant operators. The rate is notably lowest for RL submissions, in which mutants can only correct 5.61%. In contrast, LTL submissions have nearly quadrupled the fix rate at 19.74%. This trend holds for submissions with multiple fixes. Based on classification, overconstrained faults have the most fixes possible at 18.34% followed by underconstrained faults at 15.48%, while faults that have both can only be fixed 6.24% of the time. This is expected, as both over- and underconstrained faults mean the formula is sometimes too permissive and sometimes too restrictive, which is likely to lead to multi-step edits not captured by first-order mutants.
The overall low fix rate is in line with observations from RQ6 in which users over make very similar logical mistakes but the fix is actually two edit steps away, not one. For instance, there is no first order mutation operator that transforms “a in b,” into “some a & b.” In addition, there is no first-order mutant operator to transform “all a : A | some a.b” into “some A.b.” While second-order mutants would have scalability issues, the common mistakes from RQ6 could be utilized to produce a select subset that reflects common mistakes. Interestingly, mutation operators can correct some type errors, such as transforming “s.enrolled.grades” from RQ4 into “some s.enrolled.grades,” which could be utilized to improve the Analyzer’s current type-based error reports with suggested fixes.
@@
The high rate of replacement fixes is supported by observations from our common mistakes (RQ6), where we found novice users incorrectly select between similar but different operators. In particular, BOR has the second highest fix rate (19.89%), which is reflective of the fact that a vast majority of the operators that were mixed up are binary operators. QOR fixes having the third highest fix rate (17.78%) is also reflective of both choosing the wrong operator, as well as the high prevalence of FOL mistakes. Interestingly, we found that several of the QOR fixes are related to replacing one of the nested quantifiers, such as replacing “all a, some b” with “all a, one b”. For insertions, UOI fixes were dominant at 31.16% of all fixes. UOI fixes for LTL operators almost always involve inserting always, which is not unexpected as RQ4 highlights that novice users tend to push temporal operators to the front of formulas. For RL, PL, and FOL submissions, the UOI fixes were often the result of inserting transitive closure or transpose, the latter of which is interesting as it means the user was close to the right expression but had inverted part of the relational logic.
@@ 4.8. RQ8: How do novice users respond to faulty or invalid submissions?
As users incrementally work towards a correct solution, users make changes that may move their submission from one classification to the next. These trends can provide insight into how users respond to discovering their submission is faulty or invalid. To explore this, we collected all back-to-back single edit steps for an attempt. Table 13 displays the classification change, or lack thereof, produced by all of these edits. The letter E represents empty submissions. Row # Edit ‘?’ represents the number of edits that start as one classification and move to a new one based on the corresponding column, while Row % Edit ‘?’ represents what percentage this number is out of the total edits that start as classification ‘?’. For instance, there are 4,133 edits in which the user starts with an overconstrained formula and produces an overconstrained formula (O-O).
Based on Table 13, a user is significantly more likely to make an edit that does not change the classification, which implies that users often make small edits. In addition, users do frequently go from an overconstrained to a correct formula (22.92%), which is in line with our “one step away” results in which overconstrained formulas were often close to a correct submission. In contrast, syntax errors, type errors and both submissions are the most likely to lead to a student effectively resetting their attempt by erasing down to an empty submission, which indicates for these categories users are more likely to feel as if they ended up far off from the solution. Interestingly, users do go from a correct submission to all the other classifications. There are two main reasons for this. First, some of the exercises ask for an answer using a specific type of logic, but Alloy4Fun does not restrict the available operators. Therefore, users will create a correct submission and attempt to re-write it into the required logic. Second, users refine their submission, often trying to create a more condense formula.